# VD3D: Taming Large Video Diffusion Transformers for 3D Camera Control

### v2

@@ 3.2 Camera control for spatiotemporal transformers
Spatiotemporal camera representation. The standard way of representing camera parameters for a video (in the pinhole model) is via a trajectory of extrinsics and intrinsics camera parameters $(\bm{C}_{f},\bm{K}_{f})_{f=1}^{F}$ for each $f$-th frame, where $\bm{C}_{f}=[\bm{R};\bm{t}]\in\mathbb{R}^{3\times 4}$, describes the camera rotation $\bm{R}\in\mathbb{R}^{3\times 3}$ and translation $\bm{t}\times\mathbb{R}^{3}$, and $\bm{K}_{f}\in\mathbb{R}^{3\times 3}$ contains the focal length and principal point (and also horizontal/vertical skew coefficient, but it is always 0 in our setup). To control camera motion, existing methods such as [119, 135, 39] condition the temporal attention layers of U-Net-based video generators on embeddings computed from these camera parameters. Such a pipeline provides a good conditioning signal for convolutional video generators with decomposed spatial/temporal computation, but our experiments demonstrate that it works poorly for spatiotemporal transformers: they either fail to pick up any controllability (when being added as transformed residuals to the latent tokens), or degrade the visual quality of the output (when the original network parameters are being fine-tuned). This motivates us to design a better camera conditioning scheme, tailored for modern large-scale spatiotemporal transformers.
@@
First, we propose to normalize the camera parameters to the very first frame, forcing the first frame’s rotation be an identity matrix $\bm{R}_{1}=\bm{I}$ and translation be zero: $\bm{t}_{1}=\bm{0}$. We then recompute the extrinsics for each $f$-th frame as $\bm{C}_{f}=[\bm{R}_{1}^{-1}\bm{R}_{f};\bm{t}_{f}-\bm{t}_{1}]$, establishing a consistent coordinate system across videos. After that, we found it essential to enrich the conditioning information by switching from temporal frame-level camera parameters to pixel-wise spatiotemporal camera parameters. This is achieved by computing the Plücker coordinates for each pixel, providing fine-grained positional representation.
@@
Plücker coordinates provide a convenient parametrization of lines in the 3D space, and we use them to compute fine-grained positional representations of each pixel in each frame of a video. Given the extrinsic and intrinsic camera parameters $\bm{R},\bm{t}$, $\bm{K}_{f}$ of the $f$-th frame, we parametrize each $(h,w)$-th pixel as a Plücker embedding $\ddot{\bm{p}}_{f,h,w}\in\mathbb{R}^{6}$ from the camera position to the pixel’s center as
$\ddot{\bm{p}}_{f,h,w}=(\bm{t}_{f}\times\hat{\bm{d}}_{f,h,w},\hat{\bm{d}}_{f,h,w}),\quad\hat{\bm{d}}_{f,h,w}=\frac{\bm{d}}{\|\bm{d}_{f,h,w}\|},\quad\bm{d}_{f,h,w}=\bm{R}_{f}\bm{K}_{f}[w,h,1]^{\top}+\bm{t}_{f}.$ (2)
@@
This approach mirrors the technique used in recent 3D works [95, 12, 46] as well as CameraCtrl [29], a concurrent study focusing on camera control in U-Net-based video diffusion models. The motivation for using Plücker coordinates is that geometric manipulations in the Plücker space can be performed through simple arithmetics on the coordinates, which makes it easier for the network to use the positional information stored in such a disentangled representation.
@@
Computing Plücker coordinates for each pixel results in a $\ddot{\bm{P}}\in\mathbb{R}^{6\times F\times H\times W}$ spatiotemporal camera representation for a video. To input it into the model, we first perform the equivalent ViT-like [19] $h_{p}\times w_{p}$ patchification procedure. It is followed by a learnable feedforward layer $\text{FF}_{\ddot{p}}$ to obtain the Plücker camera tokens sequence $[\ddot{\bm{c}}_{\ell}]_{\ell=1}^{L}\in\mathbb{R}^{L\times d}$ of the same length $L=F\times(H/h_{p})\times(W/w_{p})$ and dimensionality $d$ as the video tokens sequence $[\bm{v}_{\ell}]_{\ell=1}^{L}$. This spatiotemporal representation carries fine-grained positional information about each pixel in a video, making it easier for the generator to accurately follow the desired camera motion.
@@
Camera conditioning. To input the rich spatiotemporal camera information in the form of Plücker embeddings into our video generator, we design an efficient ControlNet like [150] mechanism tailored for large transformer models (see Fig. 3). This mechanism is guided by two main objectives: 1) the model should be amenable to rapid fine-tuning from a small dataset with estimated camera positions; and 2) the visual quality shouldn’t be compromised during the fine-tuning stage. We found that meeting these objectives is more challenging for spatiotemporal transformers compared to U-Net-based models with decomposed spatial/temporal computation, since even minor interventions into their design quickly lead to degraded video outputs. We hypothesize that the core reason for it is the entangled spatial/temporal computation of video transformers: any attempt to alter the temporal dynamics (such as camera motion) influences spatial communication between the tokens, leading to unnecessary signal propagation and overfitting during the fine-tuning stage. To mitigate this, we input the camera information gradually through read cross-attention layers, zero-initialized from the original network parameters of the corresponding layers.
@@
Specifically, in each $b$-th FIT block of our video generator, we replace its standard read cross-attention operation (see [41, 72] for details):
$[\bm{z}^{(b)}_{m}]_{m=1}^{M}=\text{FF}^{(b)}(\text{XAttn}^{(b)}([\bm{z}^{(b)}_{m}]_{m=1}^{M},[\bm{v}^{(b)}_{\ell}]_{\ell=1}^{L}]),$ (3)
@@
where FF($\cdot$) and XAttn($\cdot$,$\cdot$ , ⋅) denote feed-forward and cross-attention layers respectively, with:
$[\bm{z}^{(b)}_{m}]_{m=1}^{M}$ $=\text{FF}^{(b)}(\text{XAttn}^{(b)}([\bm{z}^{(b)}_{m}]_{m=1}^{M},[\bm{v}^{(b)}_{\ell}]_{\ell=1}^{L}])$ (4)
[…]
@@ 3.3 Training details
Optimization details. To ensure comparability with prior work such as [119], we train our video generator on the same RealEstate10k dataset [156]. We optimize only the newly added parameters $\text{FF}_{\ddot{p}}$ and $(\alpha^{(b)},\text{FF}^{(b)},\text{XAttn}^{(b)})_{b=1}^{B}$, and keep the rest of the network frozen. We found that training only the base $36\times 64$ model was sufficient, as the $288\times 512$ upsampler already accurately follows the camera motion of a low-resolution video. We experiment with two model variants: a smaller generator with approximately 700 million parameters for ablations and initial explorations, and a larger 4 billion parameter model, which we use for the main results in this paper. Both models were trained with a batch size of 256 over 50,000 optimization steps with the LAMB optimizer [141]. The learning rate was warmed up for the first 10,000 iterations from 0 to 0.005 and then linearly decreased to 0.0015 over subsequent iterations. Since the original video diffusion model is trained in the any-frame-conditioning pipeline [72], we can produce variable camera trajectories from the same starting frame. For text conditioning, we use the T5-11B [85] language model to encode text $1024$-dimensional embeddings into 128-length sequences. For training efficiency, they were precomputed for the entire dataset. The rest of the training details have been adopted from [72].
[…]
@@ 3.4 Dataset
We fine-tune the pre-trained text-to-video model (SnapVideo [72]), on RealEstate10K [156]. The training split consists of roughly 65K video clips, and is the same as is used in concurrent work (MotionCtrl [119] and CameraCtrl [29]). We evaluate our method using 20 camera trajectories sampled from the RealEstate10K test split that were not seen during training for the user study. For the automated evaluations, we use the full test split.
@@ 3.5 Metrics
We conduct a user study to evaluate our approach for camera-controlled text-to-video generation. The study participants are presented with 20 side-by-side comparisons between the proposed approach and the baselines as well as a reference video from RealEstate10K with the same trajectory to better judge the camera alignment. We ask 20 participants for each generated video sequence to indicate which generated video they prefer based on multiple submetrics, namely, camera alignment, motion quality, text alignment, visual quality, and overall preference. The user study involved negligible risk to the participants and was conducted with appropriate institutional review board and legal approval. Participants were compensated 0.30 USD per question.
@@ 3.6 Baselines
We compare our work to the two concurrent works MotionCtrl [119] and CameraCtrl [29] by adapting their publicly released code to the SnapVideo model [72]. Note that both baselines were originally designed for space-time disentangled U-Net video diffusion models. Hence, their approaches are not directly applicable to the spatio-temporal transformers, and so we adapt them to this setting as follows. For MotionCtrl, we omit their object motion control module and use their proposed camera motion control module to encode the camera parameters into a context vector used with the SnapVideo model. We fine-tune both the camera motion control module and the cross attention between the latent vectors and the patches. We also train one version of MotionCtrl where we freeze the original base model to prevent quality degradation. For CameraCtrl, we fine-tune the original camera encoder module and use this to produce the latent vectors in the SnapVideo model. During fine-tuning the model weights are kept frozen—i.e., the same as in our proposed approach.
@@ 4.1 Assessment
We provide a qualitative and quantitative assessment of our approach compared to the baselines in Fig. 4 and in Tab. 1. Following CameraCtrl [29], we also evaluate the camera pose accuracy using ParticleSfM [152] on generated videos in Tab. 2. We using generations for text prompts from RealEstate10K [156] and MSR-VTT [129], testing both in- and out-of-distribution prompts. Note that we adjust the CameraCtrl [29] evaluation pipeline by normalizing all cameras into a unified scale as COLMAP provides different scales across different scenes. This prevents scenes with large scale to have a higher impact on the errors. Please also refer to the supplementary webpage for additional video results.
In Fig. 4 we observe that adapting the camera conditioning method from the MotionCtrl degrades visual quality and text alignment, likely because this approach adjusts the weights of the base video model. In the space-time U-Net for which this approach was proposed, the temporal layers can be fine-tuned without sacrificing visual fidelity. Since spatio-temporal transformers do not decompose temporal and spatial attributes in the same way, the model overfits to the small dataset used to fine-tune the cross-attention layer. While we observe some agreement with the camera poses used to condition the model, the text alignment is generally low in our experiments (see supplemental webpage). In contrast, CameraCtrl keeps the pre-trained video model weights frozen and only trains a camera encoder. This leads to strong visual quality, but the generated videos show little agreement with the input camera poses. We hypothesize that this is due random initialization of the new temporal attention layers and convolutions. For fair comparison, we trained all models for the same number of iterations (described in Sec. 3.3). We believe that CameraCtrl requires substantially longer training time to converge in comparison to our ControlNet-inspired conditioning mechanism (see the Plucker conditioning module in Fig. 3), which is initialized using a copy of the cross-attention and feed-forward layer and fine-tuned.
The results of the user study in Tab. 1 show that most participants prefer the generated videos using the proposed camera conditioning mechanism across all evaluated sub-metrics. We also observe a pronounced preference for the camera alignment of the proposed method compared to the other baselines. That is, 82% and 78% of participants prefer the camera alignment of the proposed method compared to our respective adaptations of MotionCtrl and CameraCtrl to the video transformer model. All results are significant at the $p<0.001$ level as evaluated using a $\chi^{2}$ test.
Figure 4: Camera-conditioned text-to-video generation. Comparison of the proposed approach to MotionCtrl and CameraCtrl for the same camera trajectory input. MotionCtrl exhibits worse quality due to fine-tuning of existing layers and CameraCtrl is not sensitive to camera conditioning. See the supplementary webpage for video results.
@@
Table 2: Camera pose evaluation. We evaluate all models using reference camera trajectories from the RealEstate10K test set. We compute translation and rotation errors based on estimated camera poses from generations using ParticleSfM [152].

### v3

@@ 3.2 Camera control for spatiotemporal transformers
Spatiotemporal camera representation. The standard way of representing camera parameters for a video (in the pinhole model) is via a trajectory of extrinsics and intrinsics camera parameters $(\bm{C}_{f},\bm{K}_{f})_{f=1}^{F}$ for each $f$-th frame. The matrix $\bm{C}_{f}=[\bm{R};\bm{t}]\in\mathbb{R}^{3\times 4}$, describes the camera rotation $\bm{R}\in\mathbb{R}^{3\times 3}$ and translation $\bm{t}\times\mathbb{R}^{3}$, and $\bm{K}_{f}\in\mathbb{R}^{3\times 3}$ contains the focal length and principal point (and also horizontal/vertical skew coefficient, but it is always 0 in our setup). To control camera motion, existing methods condition the temporal attention layers of U-Net-based video generators on embeddings computed from these camera parameters (Wang et al., 2023e; Yang et al., 2024b; Hu et al., 2024). Such a pipeline provides a good conditioning signal for convolutional video generators with decomposed spatial/temporal computation, but our experiments demonstrate that it works poorly for spatiotemporal transformers: they either fail to pick up any controllability (when being added as transformed residuals to the latent tokens), or degrade the visual quality of the output (when the original network parameters are being fine-tuned). This motivates us to design a better camera conditioning scheme, tailored for modern large-scale spatiotemporal transformers.
@@
First, we propose to normalize the camera parameters w.r.t the first frame. For this, we recompute the rotations and translations for each $f$-th frame as $\bm{R}_{f}^{\prime}=\bm{R}_{1}^{-1}\bm{R}_{f}$ and $\bm{t}_{f}^{\prime}=\bm{t}_{f}-\bm{t}_{1}$. This procedure results in normalized camera extrinsics as $\bm{C}_{f}{\color[rgb]{0,0,0}{}^{\prime}}=[\bm{R}_{f}^{\prime};\bm{t}_{f}^{\prime}]$ and establishes a consistent coordinate system across different samples in the dataset. After that, we found it essential to enrich the conditioning information by switching from temporal frame-level camera parameters to pixel-wise spatiotemporal ones. This is achieved by computing the Plücker coordinates for each pixel, providing a fine-grained positional representation.
@@
Plücker coordinates provide a convenient parametrization of lines in the 3D space, and we use them to compute fine-grained positional representations of each pixel in each frame of a video. Given the extrinsic and intrinsic camera parameters $\bm{R}{\color[rgb]{0,0,0}{}^{\prime}_{f}},\bm{t}{\color[rgb]{0,0,0}{}^{\prime}_{f}},\bm{K}_{f}$ of the $f$-th frame, we parametrize each $(h,w)$-th pixel as a Plücker embedding $\ddot{\bm{p}}_{f,h,w}\in\mathbb{R}^{6}$ from the camera position to the pixel’s center as
$\ddot{\bm{p}}_{f,h,w}=(\bm{t}_{f}{\color[rgb]{0,0,0}{}^{\prime}}\times\hat{\bm{d}}_{f,h,w},\hat{\bm{d}}_{f,h,w}),\quad\hat{\bm{d}}_{f,h,w}=\frac{\bm{d}}{\|\bm{d}_{f,h,w}\|},\quad\bm{d}_{f,h,w}=\bm{R}_{f}{\color[rgb]{0,0,0}{}^{\prime}}\bm{K}_{f}[w,h,1]^{\top}+\bm{t}_{f}{\color[rgb]{0,0,0}{}^{\prime}}.$ (2)
@@
This approach mirrors the technique used in recent 3D works (Sitzmann et al., 2021; Chen et al., 2023a; Kant et al., 2024) as well as CameraCtrl (He et al., 2024a), a concurrent study focusing on camera control in U-Net-based video diffusion models. The motivation for using Plücker coordinates is that geometric manipulations in the Plücker space can be performed through simple arithmetics on the coordinates, which makes it easier for the network to use the positional information stored in such a disentangled representation.
@@
Computing Plücker coordinates for each pixel results in a $\ddot{\bm{P}}\in\mathbb{R}^{6\times F\times H\times W}$ spatiotemporal camera representation for a video. To input it into the model, we first perform the equivalent ViT-like (Dosovitskiy et al., 2021) $h_{p}\times w_{p}$ patchification procedure. It is followed by a learnable 2-layered MLP $\text{MLP}_{\ddot{p}}$ with a GELU (Hendrycks & Gimpel, 2016) non-linearity to obtain the Plücker camera tokens sequence $\ddot{\bm{c}}_{\ell:L}\in\mathbb{R}^{L\times d}$ of the same length $L=F\times(H/h_{p})\times(W/w_{p})$ and dimensionality $d$ as the video tokens sequence $\bm{v}^{(b)}_{\ell:L}$. This spatiotemporal representation carries fine-grained positional information about each pixel in a video, making it easier for the generator to accurately follow the desired camera motion.
@@
Camera conditioning. To input the Plücker embeddings into our video generator, we design an efficient ControlNet like (Zhang et al., 2023) mechanism tailored for large transformer models (see Fig. 3). This mechanism is guided by two main objectives: 1) the model should be amenable to rapid fine-tuning from a small dataset with estimated camera positions; and 2) the visual quality shouldn’t be compromised during the fine-tuning stage. We found that meeting these objectives is more challenging for spatiotemporal transformers compared to U-Net-based models with decomposed spatial/temporal computation, since even minor interventions into their design quickly lead to degraded video outputs. We hypothesize that the core reason for it is the entangled spatial/temporal computation of video transformers: any attempt to alter the temporal dynamics (such as camera motion) influences spatial communication between the tokens, leading to unnecessary signal propagation and overfitting during the fine-tuning stage. To mitigate this, we input the camera information gradually through read cross-attention layers, zero-initialized from the original network parameters of the corresponding layers.
@@
Specifically, in each $b$-th FIT block of our video generator, we replace its standard read cross-attention operation (see (Jabri et al., 2022; Menapace et al., 2024) for details):
$\bm{z}{\color[rgb]{0,0,0}{}^{\prime}}^{(b)}_{m:M}=\text{FF}^{(b)}(\text{XAttn}^{(b)}(\bm{z}^{(b)}_{m:M},\bm{v}^{(b)}_{\ell:L})),$ (3)
@@
where FF($\cdot$) and XAttn($\cdot$,$\cdot$ , ⋅) denote feed-forward and cross-attention layers respectively. Our revised formulation is given as:
${\color[rgb]{0,0,0}\bm{z}{\color[rgb]{0,0,0}{}^{\prime}}^{(b)}_{m:M}=\text{FF}^{(b)}(\text{XAttn}^{(b)}(\bm{z}^{(b)}_{m:M},\bm{v}^{(b)}_{\ell:L}))+\text{Conv}_{\text{res}}^{(b)}\left[\text{FF}_{\text{cam}}^{(b)}(\text{XAttn}_{\text{cam}}^{(b)}(\bm{z}^{(b)}_{m:M},\ddot{\bm{c}}_{\ell:L}+\bm{v}^{(b)}_{\ell:L}))\right]},$ (4)
where $\text{FF}_{\text{cam}}^{(b)}$, and $\text{XAttn}_{\text{cam}}^{(b)}$ are learnable layers, $\text{Conv}_{\text{res}}^{(b)}$ is a $1$-dimensional convolution that processes camera-augmented latents. The produced latents $\bm{z}{\color[rgb]{0,0,0}{}^{\prime}}^{(b)}_{m:M}$ are then passed to the sequence of four self-attention layers, which are the core computational component of FIT. It is crucial to instantiate the weights of the output convolutions $\text{Conv}_{\text{res}}^{(b)}$ from zeros to preserve the model initialization. Besides, we initialize the weights of $\text{FF}_{\text{cam}}^{(b)}$ and $\text{XAttn}_{\text{cam}}^{(b)}$ from the corresponding parameters of the original network. This approach helps to preserve visual quality at initialization and facilitates rapid fine-tuning on a small dataset. As a result, we obtain the method for fine-grained 3D camera control in large video diffusion transformers. We name it VD3D and visualize its architecture in Figure 3.
@@ 3.3 Training details
To ensure comparability with prior work such as (Wang et al., 2023e), we train our video generator on the same RealEstate10K dataset (Zhou et al., 2018). We optimize only the newly added parameters $\text{FF}_{\ddot{p}}$ and $(\text{FF}^{(b)},\text{XAttn}^{(b)})_{b=1}^{B}$, and keep the rest of the network frozen. We found that training only the base $36\times 64$ model was sufficient, as the $288\times 512$ upsampler already accurately follows the camera motion of a low-resolution video.
Figure 4: Camera-conditioned text-to-video generation. Comparison of the proposed approach to MotionCtrl and CameraCtrl for the same camera trajectory input. We adapted MotionCtrl and CameraCtrl to the same transformer-based model for fair comparisons. MotionCtrl exhibits worse quality due to fine-tuning existing layers, and CameraCtrl is not sensitive to camera conditioning. See the supplementary webpage for video results.
@@ 3.4 Dataset
We fine-tune a text-to-video model, pre-trained on 2D video data, on RealEstate10K (Zhou et al., 2018). The training split for fine-tuning consists of roughly 65K video clips, and is the same as is used in concurrent work (MotionCtrl (Wang et al., 2023e) and CameraCtrl (He et al., 2024a)).
@@ 3.5 Metrics
We conduct a user study to evaluate our approach for camera-controlled text-to-video generation. The study participants are presented with 20 side-by-side comparisons between the proposed approach and the baselines as well as a reference video from RealEstate10K with the same trajectory to better judge the camera alignment. We ask 20 participants for each generated video sequence to indicate which generated video they prefer based on multiple submetrics, namely, camera alignment, motion quality, text alignment, visual quality, and overall preference. The user study involved negligible risk to the participants and was conducted with appropriate institutional review board and legal approval. We evaluate our method using 20 camera trajectories sampled from the RealEstate10K test split that were not seen during training for the user study. We use the full test split combined with unseen text prompts for the automated camera evaluations, i.e., 6928 unseen camera trajectories combined with out-of-distribution text prompts.
@@ 3.6 Baselines
We compare our work to MotionCtrl (Wang et al., 2023e) and the concurrent work CameraCtrl (He et al., 2024a) by adapting their publicly released code to the same pre-trained video model as ours. Note that both baselines were originally designed for space-time disentangled U-Net video diffusion models. Hence, their approaches are not directly applicable to the spatio-temporal transformers, and so we adapt them to this setting as follows. For MotionCtrl, we omit their object motion control module and use their proposed camera motion control module to encode the camera parameters into a context vector used with the SnapVideo model. We fine-tune both the camera motion control module and the cross attention between the latent vectors and the patches. For CameraCtrl, we fine-tune the original camera encoder module and use this to produce the latent vectors in the SnapVideo model. During fine-tuning the model weights are kept frozen—i.e., the same as in our proposed approach. Furthermore, we include comparisons to the original MotionCtrl and CameraCtrl works built on the U-Net-based AnimateDiff (Guo et al., 2024) architecture. For fair comparison, we evaluate all metrics for MotionCtrl (U-Net) and CameraCtrl (U-Net) on the same test sets as our method using the publicly provided code and checkpoints. Moreover, we provide metrics for the base model used across all our experiments, i.e., a model without camera injection. We provide more baseline variants of MotionCtrl and CameraCtrl in the appendix in Sec. B.3.
@@ 4.1 Assessment
We provide a qualitative and quantitative assessment of our approach compared to the baselines in Fig. 4 and in Tab. 1. Following CameraCtrl (He et al., 2024a), we also evaluate the camera pose accuracy using ParticleSfM (Zhao et al., 2022) on generated videos in Tab. 2. We use generations for text prompts from RealEstate10K (Zhou et al., 2018) and MSR-VTT (Xu et al., 2016), testing both in- and out-of-distribution prompts. Note that we adjust the CameraCtrl (He et al., 2024a) evaluation pipeline by normalizing all cameras into a unified scale as COLMAP provides different scales across different scenes. This prevents scenes with large scale to have a higher impact on the errors. Please also refer to the supplementary webpage for additional video results.
In Fig. 4 we observe that adapting the camera conditioning method from the MotionCtrl degrades visual quality and text alignment, likely because this approach adjusts the weights of the base video model. In the space-time U-Net for which this approach was proposed, the temporal layers can be fine-tuned without sacrificing visual fidelity. Since spatio-temporal transformers do not decompose temporal and spatial attributes in the same way, the model overfits to the small dataset used to fine-tune the cross-attention layer. While we observe some agreement with the camera poses used to condition the model, the text alignment is generally low in our experiments (see supplemental webpage). In contrast, CameraCtrl keeps the pre-trained video model weights frozen and only trains a camera encoder. This leads to strong visual quality, but the generated videos show little agreement with the input camera poses. For fair comparison, we trained all models for the same number of iterations (described in Sec. 3.3).
The results of the user study in Tab. 1 show that most participants prefer the generated videos using the proposed camera conditioning mechanism across all evaluated sub-metrics. We also observe a pronounced preference for the camera alignment of the proposed method compared to the other baselines. That is, 82% and 78% of participants prefer the camera alignment of the proposed method compared to our respective adaptations of MotionCtrl and CameraCtrl to the video transformer model. All results are significant at the $p<0.001$ level as evaluated using a $\chi^{2}$ test. We further present image-based metrics for multi-view generation and video generation quality metrics in the appendix in Sec. B.1 and Tab. B.2 respectively. The results of our camera pose accuracy evaluation in Tab. 2 further demonstrate that our approach clearly outperforms previous works. Note that the previous works MotionCtrl and CameraCtrl especially struggle with the rotation accuracy.
@@
Table 2: Camera pose evaluation. We evaluate all models using reference camera trajectories from the RealEstate10K test set. We compute translation and rotation errors based on estimated camera poses from generations using ParticleSfM (Zhao et al., 2022).